# Long distance propagation of wave beams in paraxial regime

### v1

@@ 4.1 Phase compensated wave field and preparatory calculations
$\|\mathbb{E}\psi^{\theta}(z)\|\leq\|\psi^{\theta}(0)\|e^{\frac{(p+q)^{2}R(0)z}{2\eta^{2}}}\,,$ (32)
@@
$\|\sum\limits_{\begin{subarray}{c}n,n^{\prime}=1\\ n\neq n^{\prime}\end{subarray}}^{p+q}\langle v_{n}\rangle^{\alpha^{\prime}}\langle v_{n^{\prime}}\rangle^{\alpha^{\prime}}\Big{(}\prod\limits_{m=1}^{p+q}\langle v_{m}\rangle^{\alpha}\Big{)}\mathbb{E}\psi^{\theta}(z)\|\leq\|\prod\limits_{m=1}^{p+q}\langle v_{m}\rangle^{\alpha+\alpha^{\prime}}\psi^{\theta}(0)\|e^{\frac{\lambda[R,p,q,z,\alpha]}{\eta^{2}}}\,,$ (33)
@@
$\|\mathbb{E}\psi^{\theta}(z)\|\leq\sum\limits_{N\geq 0}\frac{1}{(\eta\varepsilon\theta)^{N}}\int\limits_{[0,z]^{2N}_{<}}\int\limits_{\mathbb{R}^{Nd}}\sum\limits_{P^{2N}}\sum\limits_{j=1}^{(p+q)^{2N}}\|\psi^{\theta}(0,v-A_{j}^{\prime}\vec{k})\|\prod\limits_{(j^{\prime},l^{\prime})\in P^{2N}}\hat{\mathfrak{C}}^{\theta}(s_{j^{\prime}}-s_{l^{\prime}},k_{j^{\prime}})\frac{\mathrm{d}k_{j^{\prime}}}{(2\pi)^{d}}\prod\limits_{j=1}^{2N}\mathrm{d}s_{j}\,.$
@@
$\|\mathbb{E}\psi^{\theta}(z)\|$ $\leq\|\psi^{\theta}(0)\|\sum\limits_{N\geq 0}\frac{(p+q)^{2N}}{(\eta\varepsilon\theta)^{N}}\int\limits_{[0,z]^{2N}_{<}}\int\limits_{\mathbb{R}^{Nd}}\sum\limits_{P^{2N}}\prod\limits_{(j^{\prime},l^{\prime})\in P^{2N}}\hat{\mathfrak{C}}^{\theta}(s_{j^{\prime}}-s_{l^{\prime}},k_{j^{\prime}})\frac{\mathrm{d}k_{j^{\prime}}}{(2\pi)^{d}}\prod\limits_{j=1}^{2N}\mathrm{d}s_{j}$
$=\|\psi^{\theta}(0)\|\sum\limits_{N\geq 0}\frac{(p+q)^{2N}}{(\eta\varepsilon\theta)^{N}}\int\limits_{[0,z]^{2N}_{<}}\sum\limits_{P^{2N}}\prod\limits_{(j^{\prime},l^{\prime})\in P^{2N}}{\mathfrak{C}}^{\theta}(s_{j^{\prime}}-s_{l^{\prime}},0)\prod\limits_{j=1}^{2N}\mathrm{d}s_{j}\,.$
@@
$\|\mathbb{E}\psi^{\theta}(z)\|$ $\leq\|\psi^{\theta}(0)\|\sum\limits_{N\geq 0}\frac{1}{(2N)!}\frac{(p+q)^{2N}}{(\eta\varepsilon\theta)^{N}}\int\limits_{[0,z]^{2N}}\sum\limits_{P^{2N}}\prod\limits_{(j^{\prime},l^{\prime})\in P^{2N}}{\mathfrak{C}}^{\theta}(s_{j^{\prime}}-s_{l^{\prime}},0)\prod\limits_{j=1}^{2N}\mathrm{d}s_{j}$
$=\|\psi^{\theta}(0)\|\sum\limits_{N\geq 0}\frac{(p+q)^{2N}}{2^{N}N!}\Big{(}\frac{1}{\eta\varepsilon\theta}\int\limits_{[0,z]^{2}}\mathfrak{C}^{\theta}(s_{1}-s_{2},0)\mathrm{d}s_{1}\mathrm{d}s_{2}\Big{)}^{N}\,.$
@@
$\|\mathbb{E}\psi^{\theta}(z)\|$ $\leq\|\psi^{\theta}(0)\|\sum\limits_{N\geq 0}\frac{(p+q)^{2N}}{2^{N}N!}\Big{(}\frac{z}{\eta^{2}}R(0)\Big{)}^{N}=\|\psi^{\theta}(0)\|e^{\frac{(p+q)^{2}R(0)z}{2\eta^{2}}}\,.$
@@
×$\times\prod\limits_{(j^{\prime},l^{\prime})\in P^{2N}}\hat{\mathfrak{C}}^{\theta}(s_{j^{\prime}}-s_{l^{\prime}},k_{j^{\prime}})\frac{\mathrm{d}k_{j^{\prime}}}{(2\pi)^{d}}\prod\limits_{j=1}^{2N}\mathrm{d}s_{j}\,,$
@@
×$\times\int\limits_{[0,z]^{2N}_{<}}\int\limits_{\mathbb{R}^{Nd}}\sum\limits_{P^{2N}}\prod\limits_{(j^{\prime},l^{\prime})\in P^{2N}}\hat{\mathfrak{C}}^{\theta}(s_{j^{\prime}}-s_{l^{\prime}},k_{j^{\prime}})\langle k_{j^{\prime}}\rangle^{2+(p+q)\alpha}\frac{\mathrm{d}k_{j^{\prime}}}{(2\pi)^{d}}\prod\limits_{j=1}^{2N}\mathrm{d}s_{j}\,.$
Note that
$\int\limits_{\mathbb{R}}\int\limits_{\mathbb{R}^{d}}\langle k\rangle^{n}\hat{\mathfrak{C}}(s,k)\frac{\mathrm{d}k\mathrm{d}s}{(2\pi)^{d}}=\|\langle k\rangle^{n}\hat{R}(k)\|\,.$
@@
$\|\sum\limits_{\begin{subarray}{c}n,n^{\prime}=1\\ n\neq n^{\prime}\end{subarray}}^{p+q}\prod\limits_{m=1}^{p+q}\langle v_{m}\rangle^{\alpha_{m}}\mathbb{E}\psi^{\theta}(z)\|$ $\leq\|\prod\limits_{m=1}^{p+q}\langle v_{m}\rangle^{\alpha^{\prime}+\alpha}\psi^{\theta}(0)\|\sum\limits_{N\geq 0}\frac{C^{(p+q)N}(p+q)^{2N}}{2^{N}N!\eta^{2N}}z^{N}\|\langle k\rangle^{2+(p+q)\alpha}\hat{R}(k)\|^{N}$
$\leq\|\prod\limits_{m=1}^{p+q}\langle v_{m}\rangle^{\alpha^{\prime}+\alpha}\psi^{\theta}(0)\|e^{\frac{\lambda[R,p,q,z,\alpha]}{\eta^{2}}},$
@@
$\lambda[R,p,q,z,\alpha]=\frac{1}{2}C^{p+q}(p+q)^{2}z\|\langle k\rangle^{2+(p+q)\alpha}\hat{R}(k)\|\,.$ (36)
@@ 4.2 Convergence of moments: from paraxial to Itô-Schrödinger
$\|\prod\limits_{m=1}^{p+q}\langle v_{m}\rangle^{\alpha}\mathscr{R}^{\theta}_{p,q}(z,v)\|\leq c(p,q,z)\frac{\theta}{\eta^{2}}e^{\frac{\lambda[R,p,q,z,\alpha]}{\eta^{2}}}\|\prod\limits_{m=1}^{p+q}\langle v_{m}\rangle^{2}\psi^{\theta}(0,v)\|\,,$ (39)
where $\lambda[R,p,q,z,\alpha]=\frac{1}{2}C^{p+q}(p+q)^{2}z\|\langle k\rangle^{2+(p+q)\alpha}\hat{R}(k)\|$ for a fixed constant $C$ and $0\leq\alpha\leq\frac{d_{0}+1}{p+q}$.
@@
[…]
@@
$=-\int\limits_{\mathbb{R}^{2d}}\mathbb{E}[\hat{\nu}^{\theta}_{s_{1}}(k)\hat{\nu}^{\theta}_{s_{2}}(k^{\prime})]\Big{[}\sum\limits_{j=1}^{p}\sum\limits_{l=1}^{q}(e^{\frac{i\eta}{\varepsilon}[z_{1}g(\xi_{j},k)-z_{2}g(\zeta_{l},-k^{\prime})]}\psi(\xi_{j}-k,\zeta_{l}+k^{\prime})$
@@
$-\sum\limits_{j=1}^{p}e^{\frac{i\eta}{\varepsilon}[z_{1}g(\xi_{j},k)+z_{2}g(\xi_{j}-k,k^{\prime})]}\psi(\xi_{j}-k-k^{\prime})-\sum\limits_{l=1}^{q}e^{-\frac{i\eta}{\varepsilon}[z_{1}g(\zeta_{l},-k)+z_{2}g(\zeta_{l}+k,-k^{\prime})]}\psi(\zeta_{l}+k+k^{\prime})\Big{]}\frac{\mathrm{d}k\mathrm{d}k^{\prime}}{(2\pi)^{2d}}\,.$
Using the stationarity property of $\nu$: $\mathbb{E}[\hat{\nu}^{\theta}_{s_{1}}(k)\hat{\nu}^{\theta}_{s_{2}}(k^{\prime})]=(2\pi)^{d}\hat{\mathfrak{C}}^{\theta}(s_{1}-s_{2},k)\delta(k+k^{\prime})$, where $\mathfrak{C}^{\theta}$ is defined in (12), we deduce from the above expression that
@@
with ~$\tilde{\mathcal{B}}^{\theta}_{p,q}$ given by
~$\tilde{\mathcal{B}}^{\theta}_{p,q}(z,X,Y)=\int\limits_{\mathbb{R}^{(p+q)d}}\Pi^{\theta}(z,v)\mathcal{B}^{\theta}_{p,q}(z,v)e^{i(\sum_{j=1}^{p}\xi_{j}\cdot x_{j}-\sum_{l=1}^{q}\zeta_{l}\cdot y_{l})}\prod\limits_{j=1}^{p}\frac{\mathrm{d}\xi_{j}}{(2\pi)^{d}}\prod\limits_{l=1}^{q}\frac{\mathrm{d}\zeta_{l}}{(2\pi)^{d}}\,,$
where $\mathcal{B}^{\theta}_{p,q}$ is given by (37) and ~$\tilde{\mathscr{R}}_{p,q}^{\theta}$ satisfies
~$\|\tilde{\mathscr{R}}_{p,q}^{\theta}(z,X,Y)\|_{\infty}\leq c(p,q,z)\frac{\theta}{\eta^{2}}e^{\frac{\lambda[R,p,q,z,0]}{\eta^{2}}}\|\prod\limits_{m=1}^{p+q}\langle v_{m}\rangle^{2}\psi^{\theta}(0,v)\|\,.$
Convergence of moments: From paraxial to complex Gaussian.
@@
[…]
Control of the error terms in (39).
@@
$\|\prod\limits_{m=1}^{p+q}\langle v_{m}\rangle^{\alpha}\mathscr{R}^{\theta}_{1}(z,v)\|\leq c_{1}(p,q,z)\frac{\varepsilon\theta}{\eta^{3}}\|\prod\limits_{m=1}^{p+q}\langle v_{m}\rangle^{\alpha}\psi^{\theta}(0,v)\|e^{\frac{\lambda[R,p,q,z,\alpha]}{\eta^{2}}}\,,$ (49)
@@
$\mathbb{E}[\hat{\nu}^{\theta}_{s}(k)\hat{\nu}^{\theta}_{s^{\prime}}(k^{\prime})\prod\limits_{j=1}^{2N}\hat{\nu}^{\theta}_{s_{j}}(k_{j})]-\mathbb{E}[\hat{\nu}^{\theta}_{s}(k)\hat{\nu}^{\theta}_{s^{\prime}}(k^{\prime})]\mathbb{E}[\prod\limits_{j=1}^{2N}\hat{\nu}^{\theta}_{s_{j}}(k_{j})]$
$=\sum\limits_{\begin{subarray}{c}m,n=1\\ n\neq m\end{subarray}}^{2N}\mathbb{E}[\hat{\nu}^{\theta}_{s}(k)\hat{\nu}^{\theta}_{s_{m}}(k_{m})]\mathbb{E}[\hat{\nu}^{\theta}_{s^{\prime}}(k^{\prime})\hat{\nu}^{\theta}_{s_{n}}(k_{n})]\mathbb{E}[\prod_{\begin{subarray}{c}j=1\\ j\neq m,n\end{subarray}}^{2N}\hat{\nu}^{\theta}_{s_{j}}(k_{j})]\,,$
@@
$\mathscr{R}^{\theta}_{1}(z,v)$ $=\sum\limits_{N\geq 1}\frac{(-1)^{N+1}}{(\eta\varepsilon\theta)^{N+1}}\int\limits_{0}^{z}\int\limits_{0}^{s}\int\limits_{[0,s^{\prime}]^{2N}_{<}}\int\limits_{\mathbb{R}^{(2N+2)d}}\sum_{\begin{subarray}{c}m,n=1\\ n\neq m\end{subarray}}^{2N}\mathbb{E}[\hat{\nu}^{\theta}_{s}(k)\hat{\nu}^{\theta}_{s_{m}}(k_{m})]\mathbb{E}[\hat{\nu}^{\theta}_{s^{\prime}}(k^{\prime})\hat{\nu}^{\theta}_{s_{n}}(k_{n})]$
×$\times\mathbb{E}[\prod_{\begin{subarray}{c}j=1\\ j\neq m,n\end{subarray}}^{2N}\hat{\nu}^{\theta}_{s_{j}}(k_{j})]\sum\limits_{j=1}^{(p+q)^{2N+2}}e^{\frac{i\eta}{\varepsilon}G_{j}(\vec{s},\vec{k},v)}\psi^{\theta}(0,v-A_{j}\vec{k})\prod\limits_{j=1}^{2N+2}\frac{\mathrm{d}k_{j}}{(2\pi)^{d}}\prod\limits_{j=1}^{2N}\mathrm{d}s_{j}\mathrm{d}s^{\prime}\mathrm{d}s\,.$
@@
×^$\times\hat{\mathfrak{C}}^{\theta}(s-s_{m},k)\hat{\mathfrak{C}}^{\theta}(s^{\prime}-s_{n},k^{\prime})\frac{\mathrm{d}k\mathrm{d}k^{\prime}}{(2\pi)^{2d}}\prod\limits_{\begin{subarray}{c}(j^{\prime},l^{\prime})\in P^{2N-2}\\ j^{\prime},l^{\prime}\neq m,n\end{subarray}}\hat{\mathfrak{C}}^{\theta}(s_{j^{\prime}}-s_{l^{\prime}},k_{j^{\prime}})\frac{\mathrm{d}k_{j^{\prime}}}{(2\pi)^{d}}\prod\limits_{j=1}^{2N}\mathrm{d}s_{j}\mathrm{d}s^{\prime}\mathrm{d}s\,.$

### v2

@@ 4.1 Phase compensated wave field and preparatory calculations
$\|\mathbb{E}\psi^{\theta}(z)\|\leq\|\psi^{\theta}(0)\|e^{\frac{(p+q)^{2}z}{2\eta^{2}}{\color[rgb]{0,0,0}\int_{\mathbb{R}}\|\hat{C}(s,\cdot)\|\mathrm{d}s}}\,$, (32)
@@
$\|\sum\limits_{\begin{subarray}{c}n,n^{\prime}=1\\ n\neq n^{\prime}\end{subarray}}^{p+q}\langle v_{n}\rangle^{\alpha^{\prime}}\langle v_{n^{\prime}}\rangle^{\alpha^{\prime}}\Big{(}\prod\limits_{m=1}^{p+q}\langle v_{m}\rangle^{\alpha}\Big{)}\mathbb{E}\psi^{\theta}(z)\|\leq\|\prod\limits_{m=1}^{p+q}\langle v_{m}\rangle^{\alpha+\alpha^{\prime}}\psi^{\theta}(0)\|e^{\frac{\lambda[{\color[rgb]{0,0,0}C},p,q,z,\alpha]}{\eta^{2}}}\,$, (33)
@@
$\|\mathbb{E}\psi^{\theta}(z)\|\leq\sum\limits_{N\geq 0}\frac{1}{(\eta\varepsilon\theta)^{N}}\int\limits_{[0,z]^{2N}_{<}}\int\limits_{\mathbb{R}^{Nd}}\sum\limits_{P^{2N}}\sum\limits_{j=1}^{(p+q)^{2N}}\|\psi^{\theta}(0,v-A_{j}^{\prime}\vec{k})\|\prod\limits_{(j^{\prime},l^{\prime})\in P^{2N}}{\color[rgb]{0,0,0}|\hat{\mathfrak{C}}^{\theta}(s_{j^{\prime}}-s_{l^{\prime}},k_{j^{\prime}})|}\frac{\mathrm{d}k_{j^{\prime}}}{(2\pi)^{d}}\prod\limits_{j=1}^{2N}\mathrm{d}s_{j}\,.$
@@
$\|\mathbb{E}\psi^{\theta}(z)\|$ $\leq\|\psi^{\theta}(0)\|\sum\limits_{N\geq 0}\frac{(p+q)^{2N}}{(\eta\varepsilon\theta)^{N}}\int\limits_{[0,z]^{2N}_{<}}\int\limits_{\mathbb{R}^{Nd}}\sum\limits_{P^{2N}}\prod\limits_{(j^{\prime},l^{\prime})\in P^{2N}}|{\color[rgb]{0,0,0}\hat{\mathfrak{C}}^{\theta}(s_{j^{\prime}}-s_{l^{\prime}},k_{j^{\prime}})|}\frac{\mathrm{d}k_{j^{\prime}}}{(2\pi)^{d}}\prod\limits_{j=1}^{2N}\mathrm{d}s_{j}$
$=\|\psi^{\theta}(0)\|\sum\limits_{N\geq 0}\frac{(p+q)^{2N}}{(\eta\varepsilon\theta)^{N}}\int\limits_{[0,z]^{2N}_{<}}\sum\limits_{P^{2N}}\prod\limits_{(j^{\prime},l^{\prime})\in P^{2N}}{\color[rgb]{0,0,0}\|\hat{\mathfrak{C}}^{\theta}(s_{j^{\prime}}-s_{l^{\prime}},\cdot)\|}\prod\limits_{j=1}^{2N}\mathrm{d}s_{j}\,.$
@@
$\|\mathbb{E}\psi^{\theta}(z)\|$ $\leq\|\psi^{\theta}(0)\|\sum\limits_{N\geq 0}\frac{1}{(2N)!}\frac{(p+q)^{2N}}{(\eta\varepsilon\theta)^{N}}\int\limits_{[0,z]^{2N}}\sum\limits_{P^{2N}}\prod\limits_{(j^{\prime},l^{\prime})\in P^{2N}}{\color[rgb]{0,0,0}\|\hat{\mathfrak{C}}^{\theta}(s_{j^{\prime}}-s_{l^{\prime}},\cdot)\|}\prod\limits_{j=1}^{2N}\mathrm{d}s_{j}$
$=\|\psi^{\theta}(0)\|\sum\limits_{N\geq 0}\frac{(p+q)^{2N}}{2^{N}N!}\Big{(}\frac{1}{\eta\varepsilon\theta}\int\limits_{[0,z]^{2}}{\color[rgb]{0,0,0}\|\hat{\mathfrak{C}}^{\theta}(s_{1}-s_{2},\cdot)\|}\mathrm{d}s_{1}\mathrm{d}s_{2}\Big{)}^{N}\,.$
@@
$\|\mathbb{E}\psi^{\theta}(z)\|$ $\leq\|\psi^{\theta}(0)\|\sum\limits_{N\geq 0}\frac{(p+q)^{2N}}{2^{N}N!}\Big{(}\frac{z}{\eta^{2}}{\color[rgb]{0,0,0}\int_{\mathbb{R}}\|\hat{C}(s,\cdot)\|\mathrm{d}s}\Big{)}^{N}=\|\psi^{\theta}(0)\|e^{\frac{(p+q)^{2}z}{2\eta^{2}}{\color[rgb]{0,0,0}\int_{\mathbb{R}}\|\hat{C}(s,\cdot)\|\mathrm{d}s}}\,.$
@@
$\times\prod\limits_{(j^{\prime},l^{\prime})\in P^{2N}}{\color[rgb]{0,0,0}|\hat{\mathfrak{C}}^{\theta}(s_{j^{\prime}}-s_{l^{\prime}},k_{j^{\prime}})|}\frac{\mathrm{d}k_{j^{\prime}}}{(2\pi)^{d}}\prod\limits_{j=1}^{2N}\mathrm{d}s_{j}\,$,
@@
$\times\int\limits_{[0,z]^{2N}_{<}}\int\limits_{\mathbb{R}^{Nd}}\sum\limits_{P^{2N}}\prod\limits_{(j^{\prime},l^{\prime})\in P^{2N}}{\color[rgb]{0,0,0}|\hat{\mathfrak{C}}^{\theta}(s_{j^{\prime}}-s_{l^{\prime}},k_{j^{\prime}})|}\langle k_{j^{\prime}}\rangle^{2+(p+q)\alpha}\frac{\mathrm{d}k_{j^{\prime}}}{(2\pi)^{d}}\prod\limits_{j=1}^{2N}\mathrm{d}s_{j}\,.$
@@
$\|\sum\limits_{\begin{subarray}{c}n,n^{\prime}=1\\ n\neq n^{\prime}\end{subarray}}^{p+q}\prod\limits_{m=1}^{p+q}\langle v_{m}\rangle^{\alpha_{m}}\mathbb{E}\psi^{\theta}(z)\|$
$\leq\|\prod\limits_{m=1}^{p+q}\langle v_{m}\rangle^{\alpha^{\prime}+\alpha}\psi^{\theta}(0)\|\sum\limits_{N\geq 0}\frac{C^{(p+q)N}(p+q)^{2N}}{2^{N}N!\eta^{2N}}z^{N}{\color[rgb]{0,0,0}\Big{(}\int_{\mathbb{R}^{d+1}}\langle k\rangle^{2+(p+q)\alpha}|\hat{C}(s,k)|\mathrm{d}s\mathrm{d}k\Big{)}}^{N}$
$\leq\|\prod\limits_{m=1}^{p+q}\langle v_{m}\rangle^{\alpha^{\prime}+\alpha}\psi^{\theta}(0)\|e^{\frac{\lambda[{\color[rgb]{0,0,0}C},p,q,z,\alpha]}{\eta^{2}}},$
@@
$\lambda[{\color[rgb]{0,0,0}C},p,q,z,\alpha]=\frac{1}{2}C^{p+q}(p+q)^{2}z{\color[rgb]{0,0,0}\int_{\mathbb{R}^{d+1}}\langle s\rangle\langle k\rangle^{2+(p+q)\alpha}|\hat{C}(s,k)|\mathrm{d}s\mathrm{d}k}\,.$ (36)
@@ 4.2 Convergence of moments: from paraxial to Itô-Schrödinger
$\|\prod\limits_{m=1}^{p+q}\langle v_{m}\rangle^{\alpha}\mathscr{R}^{\theta}_{p,q}(z,v)\|\leq c(p,q,z)\frac{\theta}{\eta^{2}}e^{\frac{\lambda[{\color[rgb]{0,0,0}C},p,q,z,\alpha]}{\eta^{2}}}\|\prod\limits_{m=1}^{p+q}\langle v_{m}\rangle^{2}\psi^{\theta}(0,v)\|\,$, (39)
where $\lambda[{\color[rgb]{0,0,0}C},p,q,z,\alpha]=\frac{1}{2}C^{p+q}(p+q)^{2}z{\color[rgb]{0,0,0}\int_{\mathbb{R}^{d+1}}\langle s\rangle\langle k\rangle^{2+(p+q)\alpha}|\hat{C}(s,k)|\mathrm{d}s\mathrm{d}k}$ for a fixed constant $C$ and $0\leq\alpha\leq\frac{d_{0}+1}{p+q}$.
@@
$\int_{0}^{z}\int_{0}^{s}a(s,s^{\prime})ds^{\prime}ds$ $=\int_{0}^{z}\int_{s}^{z}a(s^{\prime},s)ds^{\prime}ds=\int_{0}^{z}\Big{(}\int_{s}^{\infty}-\int_{z}^{\infty}\Big{)}a(s^{\prime},s)ds^{\prime}ds$
@@
$=-\int\limits_{\mathbb{R}^{2d}}{\color[rgb]{0,0,0}\frac{\mathbb{E}[\hat{\nu}^{\theta}_{s_{1}}(\mathrm{d}k)\hat{\nu}^{\theta}_{s_{2}}(\mathrm{d}k^{\prime})]}{(2\pi)^{2d}}}\Big{[}\sum\limits_{j=1}^{p}\sum\limits_{l=1}^{q}(e^{\frac{i\eta}{\varepsilon}[z_{1}g(\xi_{j},k)-z_{2}g(\zeta_{l},-k^{\prime})]}\psi(\xi_{j}-k,\zeta_{l}+k^{\prime})$
@@
$-\sum\limits_{j=1}^{p}e^{\frac{i\eta}{\varepsilon}[z_{1}g(\xi_{j},k)+z_{2}g(\xi_{j}-k,k^{\prime})]}\psi(\xi_{j}-k-k^{\prime})-\sum\limits_{l=1}^{q}e^{-\frac{i\eta}{\varepsilon}[z_{1}g(\zeta_{l},-k)+z_{2}g(\zeta_{l}+k,-k^{\prime})]}\psi(\zeta_{l}+k+k^{\prime})\Big{]}\,.$
Using the stationarity property of $\nu$: $\color[rgb]{0,0,0}{\mathbb{E}[\hat{\nu}^{\theta}_{s_{1}}(\mathrm{d}k)\hat{\nu}^{\theta}_{s_{2}}(\mathrm{d}k^{\prime})]=(2\pi)^{d}\hat{\mathfrak{C}}^{\theta}(s_{1}-s_{2},k)\delta(k+k^{\prime})\mathrm{d}k}$, where $\mathfrak{C}^{\theta}$ is defined in (12), we deduce from the above expression that
@@
with $\tilde{\mathcal{B}}^{\theta}_{p,q}$ given by
$\tilde{\mathcal{B}}^{\theta}_{p,q}(z,X,Y)=\int\limits_{\mathbb{R}^{(p+q)d}}\Pi^{\theta}(z,v)\mathcal{B}^{\theta}_{p,q}(z,v)e^{i(\sum_{j=1}^{p}\xi_{j}\cdot x_{j}-\sum_{l=1}^{q}\zeta_{l}\cdot y_{l})}\prod\limits_{j=1}^{p}\frac{\mathrm{d}\xi_{j}}{(2\pi)^{d}}\prod\limits_{l=1}^{q}\frac{\mathrm{d}\zeta_{l}}{(2\pi)^{d}}\,$,
where $\mathcal{B}^{\theta}_{p,q}$ is given by (37) and $\tilde{\mathscr{R}}_{p,q}^{\theta}$ satisfies
$\|\tilde{\mathscr{R}}_{p,q}^{\theta}(z,X,Y)\|_{\infty}\leq c(p,q,z)\frac{\theta}{\eta^{2}}e^{\frac{\lambda[{\color[rgb]{0,0,0}C},p,q,z,0]}{\eta^{2}}}\|\prod\limits_{m=1}^{p+q}\langle v_{m}\rangle^{2}\psi^{\theta}(0,v)\|\,.$
Convergence of moments: From paraxial to complex Gaussian
@@
It thus remains to conclude the proof of Theorem 4.3 by proving the error bound (39). We first make the following preparations. Recall that $\|\cdot\|$ denotes the total variation norm.
Control of the error terms in (39)
@@
$\|\prod\limits_{m=1}^{p+q}\langle v_{m}\rangle^{\alpha}\mathscr{R}^{\theta}_{1}(z,v)\|\leq c_{1}(p,q,z)\frac{\varepsilon\theta}{\eta^{3}}\|\prod\limits_{m=1}^{p+q}\langle v_{m}\rangle^{\alpha}\psi^{\theta}(0,v)\|e^{\frac{\lambda[{\color[rgb]{0,0,0}C},p,q,z,\alpha]}{\eta^{2}}}\,$, (49)
@@
$\mathbb{E}[\hat{\nu}^{\theta}_{s}(\mathrm{d}k)\hat{\nu}^{\theta}_{s^{\prime}}(\mathrm{d}k^{\prime})\prod\limits_{j=1}^{2N}\hat{\nu}^{\theta}_{s_{j}}(\mathrm{d}k_{j})]-\mathbb{E}[\hat{\nu}^{\theta}_{s}(\mathrm{d}k)\hat{\nu}^{\theta}_{s^{\prime}}(\mathrm{d}k^{\prime})]\mathbb{E}[\prod\limits_{j=1}^{2N}\hat{\nu}^{\theta}_{s_{j}}(\mathrm{d}k_{j})]$
$=\sum\limits_{\begin{subarray}{c}m,n=1\\ n\neq m\end{subarray}}^{2N}\mathbb{E}[\hat{\nu}^{\theta}_{s}(\mathrm{d}k)\hat{\nu}^{\theta}_{s_{m}}(\mathrm{d}k_{m})]\mathbb{E}[\hat{\nu}^{\theta}_{s^{\prime}}(\mathrm{d}k^{\prime})\hat{\nu}^{\theta}_{s_{n}}(\mathrm{d}k_{n})]\mathbb{E}[\prod_{\begin{subarray}{c}j=1\\ j\neq m,n\end{subarray}}^{2N}\hat{\nu}^{\theta}_{s_{j}}(\mathrm{d}k_{j})]\,$,
@@
$\mathscr{R}^{\theta}_{1}(z,v)$ $=\sum\limits_{N\geq 1}\frac{1}{(\eta\varepsilon\theta)^{N+1}}\int\limits_{0}^{z}\int\limits_{0}^{s}\int\limits_{[0,s^{\prime}]^{2N}_{<}}\int\limits_{\mathbb{R}^{(2N+2)d}}\sum_{\begin{subarray}{c}m,n=1\\ n\neq m\end{subarray}}^{2N}{\color[rgb]{0,0,0}\mathbb{E}[\hat{\nu}^{\theta}_{s}(\mathrm{d}k)\hat{\nu}^{\theta}_{s_{m}}(\mathrm{d}k_{m})]\mathbb{E}[\hat{\nu}^{\theta}_{s^{\prime}}(\mathrm{d}k^{\prime})\hat{\nu}^{\theta}_{s_{n}}(\mathrm{d}k_{n})]}$
$\times{\color[rgb]{0,0,0}\mathbb{E}[\prod_{\begin{subarray}{c}j=1\\ j\neq m,n\end{subarray}}^{2N}\hat{\nu}^{\theta}_{s_{j}}(\mathrm{d}k_{j})]}\sum\limits_{j=1}^{(p+q)^{2N+2}}e^{\frac{i\eta}{\varepsilon}G_{j}(\vec{s},\vec{k},v)}\psi^{\theta}(0,v-A_{j}\vec{k})\frac{\mathrm{d}s^{\prime}\mathrm{d}s}{(2\pi)^{2d}}\prod\limits_{j=1}^{2N}\frac{\mathrm{d}s_{j}}{(2\pi)^{d}}\,.$
@@
$\times{\color[rgb]{0,0,0}|\hat{\mathfrak{C}}^{\theta}(s-s_{m},k)||\hat{\mathfrak{C}}^{\theta}(s^{\prime}-s_{n},k^{\prime})|}\frac{\mathrm{d}k\mathrm{d}k^{\prime}}{(2\pi)^{2d}}\prod\limits_{\begin{subarray}{c}(j^{\prime},l^{\prime})\in P^{2N-2}\\ j^{\prime},l^{\prime}\neq m,n\end{subarray}}{\color[rgb]{0,0,0}|\hat{\mathfrak{C}}^{\theta}(s_{j^{\prime}}-s_{l^{\prime}},k_{j^{\prime}})|}\frac{\mathrm{d}k_{j^{\prime}}}{(2\pi)^{d}}\prod\limits_{j=1}^{2N}\mathrm{d}s_{j}\mathrm{d}s^{\prime}\mathrm{d}s\,.$